OpenStudy (hocuspucus292333393):

CAN u help me plz What is the quotient? −5 1/4 ÷2 3/4 Enter your answer as a mixed number, in simplified form, in the box.

OpenStudy (hocuspucus292333393):

@Jerry45 help me

OpenStudy (hocuspucus292333393):

@zoekeya here

OpenStudy (zoekeya):

I will still think. Help? @brucebaner

OpenStudy (hocuspucus292333393):

is it 1 10/11

OpenStudy (hocuspucus292333393):

am I right or what

OpenStudy (phi):

First change - 5 ¼ into an improper fraction do you know how to do that ?

OpenStudy (phi):

yes but it's negative \[ - \frac{21}{4} \] next change 2 ¾ into an improper fraction what do you get ?

OpenStudy (hocuspucus292333393):

11/4

OpenStudy (phi):

yes. that means the problem is the same as \[ -\frac{21}{4} \div \frac{11}{4} \] Now we use the idea that dividing is the same as "flipping" and then multiplying There is a way to show why this works, but for now, let's just use that idea In other words, the problem becomes \[ -\frac{21}{4} \times \frac{4}{11} \]

OpenStudy (hocuspucus292333393):

-21/4 x 4/11= 21/11

OpenStudy (phi):

when doing \[ -\frac{21}{4} \times \frac{4}{11} \] we would multiply top times top and bottom times bottom But, it's always easier to first divide (if you can do it evenly) in this problem we have a 4 "up top" and 4 "in the bottom" 4 divided by 4 is 1, and we can ignore multiply by 1 in other words, divide the top 4 by 4 (to get 1) and divide the bottom 4 by 4 to get 1 \[ -\frac{21}{\cancel{4}} \times \frac{\cancel{4}}{11} \] 11 does not divide 11 into 21, so we are done simplifying multiply top times top and bottom times bottom we get \[ - \frac{21}{11}\] (remember it's negative because we are doing a minus times a plus)

OpenStudy (phi):

we are not yet finished. they want a mixed number. can you change -21/11 into a mixed number ?

OpenStudy (hocuspucus292333393):

I'm doing it alright

OpenStudy (phi):

and remember the answer will be negative

OpenStudy (hocuspucus292333393):

-1 10/11

OpenStudy (phi):

good job

OpenStudy (hocuspucus292333393):

so that's the answer